Calculate the minimum resistance to limit the gate current (in the worst case) to 70.71% IGM on a 2N6508 SCR that is on a 100V RMS / 50Hz line. (Use a protection diode on the gate circuit. )
Answer #1

According to the datasheet of 2N6058 SCR maximum value of IGM is 2 Amp and we have to limit it to the 70.71%.

So the maximum current would be 2*0.7071=1.4142 Amp

Now the minimum resistance would be = (V/I)

R=Vrms/I =100/1.4142 = 70.71 ohm.

I u have any doubt regarding solution then please let me know in comments.